J Dexter used his savings to upgrade his sound system business. He spent 1/4 of his savings on new speakers, 1/6 of his savings on new laser lights, 1/10 of his savings on a new mixer, and 1/12 of his savings to upgrade his office. what fraction of his savings was left


1
Expert's answer
2022-02-17T15:27:02-0500

Savings left = 1-{1/4+1/6+1/10+1/12}

Savings left = 1-36/60

Savings left = 24/60

Savings left = 2/5